The reaction shown below has an equilibrium constant value of K, 8.84 at a certain temperature. 2 NO (g) N2 (g) + O2 (g) Kc = 8.84 0 moles of NO are sealed in a 2.00 L container. Calculate the concentration of Nz in the container when equilibrium is reached
media%2F538%2F5388117a-3847-4549-a9e1-0a
Answer #1

Solution

i)                            2NO (g)                =             N2 (g)     +             O2 (g)

Initial                     0.500                                     0                              0

At equilibrium (0.5 –2x)/2 x/2 x/2

Kc = [N2] [O2] / [NO]2

8.84 = (x/2)2/[(0.5 –2x)/2]2

Or 34.36x2-17.6x+2.21=0

x = 0.214

The concentration of N2 (g) at equilibrium, [N2]= x/2 = 0.214/2 = 0.107 mole/L

ii)

                                CO (g)                 +             Cl2 (g)     =                             COCl2 (g)            

At equilibrium 0.5 atm 0.225 atm                                    x

Kp = PCOCl2 / PCO x PCl2

5.5 = x / (0.5)(0.225)

x = 0.619 atm

The pressure of COCl2 at equilibrium is 0.619 atm.